OpenStudy (anonymous):

The side lengths of a cube is a. Describe the pieces that need to be added in order to increase its length, width, and height by 0.3 unit each

OpenStudy (anonymous):

\[\left(a+\frac{3}{10}\right)^3=a^3+\frac{9 a^2}{10}+\frac{27 a}{100}+\frac{27}{1000} \]\[\frac{9 a^2}{10}+\frac{27 a}{100}+\frac{27}{1000} \]